Elorus:

F(g) = 1100(1.10)x X = 20

Tranquility:

Are you trying to find the value of F(g) at x=20? Just replace x=20 and simplify \( F(g) = 1100(1.10)^x\) \( F(g) = 1100(1.10)^{20}\) You can use a calculator to first find what 1.10^20 is and then multiply it by 1100